# Dynamical tide modified Roche limit in eccentric, asynchronous binaries

### v1

@@ 2 Nonlinear hydrodynamics
For the Roche limit calculation, it is sufficient to consider only the quadrupolar ($l=2$) fundamental modes (f-modes). We drop the $l\geq 3$ perturbations, restrict our analysis to aligned spins when rotation is incorporated, and ignore the inertial modes (W. C. G. Ho & D. Lai, 1999; D. Lai & Y. Wu, 2006; H. J. Braviner & G. I. Ogilvie, 2015; W. Xu & D. Lai, 2017). Therefore, we label the modes only with their azimuthal quantum numbers $m$. The eigenfunctions $\boldsymbol{\xi}_{m}$ are computed from a non-rotating background as (e.g., E. Poisson & C. M. Will 2014)
@@
with $\epsilon=M^{\prime}/D^{3}$. We use $D$ and $\Phi$ to denote orbital separation and phase. The derivation follows a variational principle and is presented in detail in Appendix A. Note that in a homogeneous ellipsoid, the perturbation satisfies the affine model of B. Carter & J. P. Luminet (1983, 1985); P. Diener et al. (1995), where the perturbed position $\boldsymbol{r}$ is related to the unperturbed position $\boldsymbol{x}$ via
@@
where the mapping between $q_{ij}(t)$ and $q_{m}(t)$ is given in Eq. (A2). The affine model allows the exact nonlinear form of $K_{m}$ to be obtained (available in Appendix A, Eqs. A25-A30, and allows for arbitrary values of $n$ and $\Gamma$ with the caveat that Eq. 12 may not be accurate beyond $n=1/(\Gamma-1)=0$). Nonetheless, we expand the results to the three-wave order (that is, we count $q_{m}\sim\epsilon$ and include terms to $\mathcal{O}(\epsilon^{3})$ in the energy variations and $\mathcal{O}(\epsilon^{2})$ in the equation of motion). It is already sufficient to reproduce the Roche limit at this order. More importantly, such expanded equations can also be obtained following T. Van Hoolst (1994), which takes a general set of linear eigenmodes as input to compute their nonlinear couplings and applies to generic compressible fluids (realistic stars and planets). This is advantageous over the affine model, which only self-consistently describes idealized homogeneous ellipsoids. Therefore, our study bridges the two different formulations of nonlinear hydrodynamics. The exact agreement between the two formulations at the three-wave order for homogeneous ellipsoids is presented in Appendix B.
@@
with $v_{m}=(q_{m}e^{im\phi})_{\phi=\phi_{q}}$. Treating $(-v_{0})\sim v_{2}\sim\xi$, we see that Eq. (9) takes the same form as the toy model of Eq. (1), and the nonlinear terms in $K_{m}$ tend to destabilize the system when $\xi>0$. Thus, instability is most likely to occur when $\xi$ reaches its maximum along the major axis. In the equilibrium limit (without mode resonances; following the convention of H. Yu et al. 2024), where the tidal bulge follows the companion, we have $\phi_{q}=\Phi$, matching the intuition that mass transfer first occurs at the L1 Lagrangian point. At the quadrupolar order, the L2 point ($\phi=\Phi+\pi$) destabilizes at the same time.
@@
The second correction comes from the centrifugal force whose detailed incorporation is presented in Appendix A. Since the force can be derived from a specific centrifugal potential $d\mathcal{R}/dM=-\omega^{2}(r^{2}-z^{2})/2$, its impact on the equations of motion can be obtained through a variational principle by considering the energy $\mathcal{R}$. When doing so, it is important to introduce an $l=0$ mode whose amplitude we denote as $q_{\rm tr}$ (a perturbation of the trace of $q_{ij}$ in P. Diener et al. 1995) and allow it to be coupled to the $q_{m}$ mode at the three-wave level to properly capture the correction at $\mathcal{O}(\epsilon^{2})$. The same result can also be obtained from the formulation of T. Van Hoolst (1994) by splitting $d\mathcal{R}/dM=(d\mathcal{R}/dM)_{l=0}+(d\mathcal{R}/dM)_{l=2}=-\omega^{2}r^{2}/3-\omega^{2}(x^{2}+y^{2}-2z^{2})/6$, and adding $(d\mathcal{R}/dM)_{l=0}$ to the background (monopolar) gravitational potential and $(d\mathcal{R}/dM)_{l=2}$ to the first-order, quadrupolar Eulerian potential perturbation (same as the tidal interaction potential). Note that our approach treats the centrifugal potential as a (nonlinear) drive of the modes defined by a non-rotating background, instead of redefining a new set of eigenmodes of a non-spherical background where the centrifugal perturbation is absorbed. In other words, the centrifugal potential is treated on an equal footing as the tidal potential.
@@
This splitting is based on the phase-space expansion of A. K. Schenk et al. (2002), and the Coriolis force is absorbed as a frequency shift (D. Lai, 2021; H. Yu et al., 2024; H. Yu & S. Y. Lau, 2025),
@@
Higher-order corrections from the Coriolis force are ignored (but see, e.g., J. W. Dewberry & D. Lai 2022). The corresponding second-order equation can be obtained by differentiating $\dot{c}_{m,\pm}$ one more time and eliminating the first-order derivatives with the equations of motion, leading to
@@
Summarizing, the equations of motion we will solve in §2.1 and §2.2 are
@@
where $\bar{K}_{m}=K_{m}e^{im\phi_{q}}$ and $\bar{C}_{m}^{\rm(st)}=C_{m}^{\rm(st)}e^{im\phi_{q}}$. In the equilibrium limit (as considered in §2.1 and 2.2) without free oscillations of the mode due to resonances, $\phi_{a}=\Phi_{C}$, and one just needs to replace $q_{m}$ with $v_{m}$ and drop the phases following $\epsilon$ in $K_{m}$ to obtain $\bar{K}_{m}$. Eq. (22) is real valued.
2.1 Consistency with previous analyses
Figure 1: Equilibrium sequences under the static limit. Curves in gray correspond to synchronized donors (with $p=1$), while those in yellow are for the non-rotating case ($p=0$). The solid and dashed lines are evaluated at the three-wave and four-wave orders, respectively. Vertical lines mark the locations where $\epsilon$ reaches the maxima. The classic results of S. Chandrasekhar (1963) are shown in the red horizontal lines.
First, we consider the case where $M$ moves in a circular orbit with synchronous rotation, so the left-hand sides of Eq. (19) or (22) vanish. Our goal is to establish the consistency between our formulation based on an order-by-order expansion with the classic analyses of S. Chandrasekhar (1963) under this static limit. We write the spin frequency $\omega$ as $\omega^{2}=p\epsilon$, with $p=1+M/M^{\prime}$ for synchronized rotation in a circular orbit. Note that our $p$ differs from the one in S. Chandrasekhar (1963) by 1, as $p=p^{\rm(Ch)}+1$. The superscript “(Ch)” will be used when we quote notations or results from S. Chandrasekhar (1963). In our analysis, we will focus on $M^{\prime}\gg M$ (as for planet migration or TDEs), so a synchronous donor has $p=1$.
Inspired by S. Chandrasekhar (1963), we note that the perturbed fluid can be described by an ellipsoid with $\bar{x}^{2}/a_{1}^{2}+\bar{y}^{2}/a_{2}^{2}+\bar{z}^{2}/a_{3}^{2}=1$ at the surface. The barred coordinates are aligned with the axes of the ellipsoid. In the limit with circular orbit and synchronous rotation, the barred frame coincides with the corotating frame. The axes are given by (Eqs. A5 and A31)
@@ 2.1 Consistency with previous analyses
If there exists a $v_{2}$ that simultaneously solves the above two polynomials, then the resultant of the two polynomials must vanish (E. W. Weisstein, 2025). That is,
@@
In our calculation, we find that it is sufficient to capture the classic results of S. Chandrasekhar (1963) with percent-level accuracy using only three-wave interactions (solid lines; Eqs. 20 and 21). For the synchronous ($p=1$) case, we find $\epsilon_{\rm max}\simeq 0.0747$ at $\cos^{-1}a_{31}\simeq 64.8^{\circ}$, meaning the threshold separation is at $D^{(\rm th)}=(1/\epsilon_{\rm max})^{1/3}R_{t}\simeq 2.38R_{t}$, matching the classic Roche limit (indicated by the horizontal dotted lines) of $2.44R_{t}$ to $2.5\%$. A similar level of accuracy is also obtained for the non-rotating $p=0$ case, with $\epsilon_{\rm max}\simeq 0.110$, $(1/\epsilon_{\rm max})^{1/3}\simeq 2.09$, achieved at $\cos^{-1}a_{31}\simeq 65.4^{\circ}$.
The accuracy can be improved further by including the next order, four-wave interactions. The equations of motion at this order are derived in Appendix A, Eqs. (A33a) and (A33b). Still using Eq. (24) to eliminate $v_{0}$ in terms of $v_{2}$ and $a_{31}$,111 Strictly, at the four-wave order one should account for corrections to $a_{1,2,3}$ from the monopolar mode, $q_{\rm tr}$; Eqs. (A5) and (A31). This introduces nonlinear corrections to Eq. (24), which we ignore here. we again obtain two polynomial equations of $v_{2}$,
@@
by setting $\ddot{q}_{m}=0$, with $m=2$ and 0. The resultant (now functions of $W_{mn}$) needs to vanish as before for the two equations to hold simultaneously. We can thus find the equilibrium sequences at the four-wave order, shown in the dashed lines in Fig. 1. The result now matches those from S. Chandrasekhar (1963) almost exactly. For the $p=1$ case shown in gray, we have $\epsilon_{\rm max}\simeq 0.0687$ at $\cos^{-1}a_{31}=61.1^{\circ}$, corresponding to a Roche limit of $D^{(\rm th)}=(1/\epsilon_{\rm max})^{1/3}R_{t}=2.44R_{t}$.
@@
While the main result of this Subsection is to reproduce the classic ones by S. Chandrasekhar (1963) on homogeneous ellipsoids, it is significant for two reasons. First, it validates our formulation based on which $\ddot{q}_{m}\neq 0$ corrections will be introduced. Second, it bridges the analyses of S. Chandrasekhar (1963) on idealized ellipsoids and T. Van Hoolst (1994) on general stars and planets (see also N. N. Weinberg et al. 2012; N. N. Weinberg 2016; H. Yu et al. 2023). In general, exact nonlinear expressions are challenging to find, but an order-by-order expansion as in T. Van Hoolst (1994) is straightforward to compute (though lengthy). Our result provides a framework for how one can start from T. Van Hoolst (1994) to determine the Roche limit of a realistic star/planet theoretically.
2.2 Finite-frequency corrections in asynchronous, low eccentricity systems
@@
In this Subsection, we will extend the stability analysis to analytically consider finite-frequency corrections due to the equilibrium tide (following the convention of H. Yu et al. 2024), where the perturbation varies with the instantaneous tidal forcing. In other words, the tidal bulge still points towards the companion. The component that varies at a mode’s natural frequency and depends on the excitation history of the mode (what we call the dynamical tide) will be deferred to the numerical investigation in §2.3 and ignored in the analysis here. This means that the prediction derived here will be accurate for binaries with low eccentricity $e\lesssim 0.3$. For more eccentric binaries, the dynamical tides can exhibit complicated, often chaotic, evolutions that may trigger instabilities at separations greater than the prediction of this Subsection by tens of percent.
@@
To proceed, we write $q_{m}=(1+z_{m})q_{m}^{(0)}$ where $q_{m}^{(0)}\simeq I_{m}\epsilon e^{-im\Phi_{C}}$ represents an approximate solution of the perturbation, with $I_{2}=3/4$ and $I_{0}=-\sqrt{6}/4$ (Eqs. 20 and 21). We treat $z_{m}$ as a low-frequency variable (whose time derivatives will be ignored) that we will use to determine the equilibrium structure of the donor. Note Eq. (22) still holds as long as we incorporate the time derivatives on the left-hand side due to $\ddot{q}_{m}\simeq\ddot{q}_{m}^{(0)}$. In essence, this model corresponds to the toy model described in Eq. (5).
@@
The procedure for finding the Roche limit follows identically those described in §2.1. Choosing again a linear relation between $v_{0}$ and $v_{2}$ as in Eq. (24), Eq. (22) now becomes two quadratic equations of $z_{2}$:
@@
For practical usefulness, we show in Fig. 2 the threshold pericenter separation, $D_{p}^{(\rm th)}/R_{t}=(1/\epsilon_{\rm max})^{1/3}$, below which hydrodynamical instability can occur. Different colors represent different spin rates relative to the pseudo-synchronization values $\omega_{\rm ps}$ in P. Hut (1981),
@@
We pause here to discuss further the meaning of Fig. 2. As the coefficients in Eq. (29) are evaluated at the pericenter, Fig. 2 should be interpreted as the instantaneous instability threshold at the pericenter. Therefore, a donor $M$ marginally unstable at the pericenter may become stable once it moves away from the pericenter and $\epsilon$ (defined by the instantaneous separation) drops below the instability threshold. As a result, the growth of hydrodynamical instability during a single pericenter passage may not disrupt $M$ sufficiently to cause mass loss. Nonetheless, as long as the fluid dissipation timescale is long compared to the orbital period (which is very well satisfied for the large-scale f-mode when 1e>0.011-e>0.01 under the linear theory assuming damping due to turbulent convection; P. Goldreich & P. D. Nicholson 1977; Y. Wu 2005), the growth of instability at every pericenter passage will accumulate over multiple orbits, eventually disrupting $M$ and triggering mass loss (see also numerical examples to be presented in the next Subsection). For systems with low eccentricities $e\lesssim 0.3$, Fig. 2 provides a decent approximation of the onset of mass loss as a runaway in mode amplitude occurs within a few (<10<10) orbits. For systems with higher eccentricities, the amount of growth per passage is smaller, given the smaller fraction of time $M$ spends near the pericenter ($2\pi/(\dot{\Phi}_{p}P_{\rm orb})\propto\sqrt{(1-e)^{3}/(1+e)}$). This requires more orbits for the perturbation to build up until a runaway eventually occurs. However, the dynamical tide component can also be excited in those systems, significantly lowering the threshold $\epsilon$ for a stable configuration (see Eq. 5 and note the dynamical tide plays the role of $\xi_{\rm dyn}$ in the toy model). In the weak damping case, mass loss from $M$ can occur at $D_{p}$ significantly greater than the prediction of Fig. 2 for highly eccentric systems.
2.3 Dynamical tide induced instability and chaos

### v2

@@ II Nonlinear hydrodynamics
For the Roche limit calculation, it is sufficient to consider only the quadrupolar ($l=2$) fundamental modes (f-modes). We drop the $l\geq 3$ perturbations, restrict our analysis to aligned spins when rotation is incorporated, and ignore the inertial modes (Ho and Lai, 1999; Lai and Wu, 2006; Braviner and Ogilvie, 2015; Xu and Lai, 2017). Therefore, we label the modes only with their azimuthal quantum numbers $m$. The eigenfunctions $\boldsymbol{\xi}_{m}$ are computed from a non-rotating background as (e.g., Poisson and Will 2014)
@@
with $\epsilon=M^{\prime}/D^{3}$. We use $D$ and $\Phi$ to denote orbital separation and phase. The derivation follows a variational principle and is presented in detail in Appendix A. Note that in a homogeneous ellipsoid, the perturbation satisfies the affine model of Carter and Luminet (1983, 1985); Diener et al. (1995), where the perturbed position $\boldsymbol{r}$ is related to the unperturbed position $\boldsymbol{x}$ via
@@
where the mapping between $q_{ij}(t)$ and $q_{m}(t)$ is given in Eq. (A2). The affine model allows the exact nonlinear form of $K_{m}$ to be obtained (available in Appendix A, Eqs. A25-A30, and allows for arbitrary values of $n$ and $\Gamma$ with the caveat that Eq. 13 may not be accurate beyond $n=1/(\Gamma-1)=0$). Nonetheless, we expand the results to the three-wave order (that is, we count $q_{m}\sim\epsilon$ and include terms to $\mathcal{O}(\epsilon^{3})$ in the energy variations and $\mathcal{O}(\epsilon^{2})$ in the equation of motion). It is already sufficient to reproduce the Roche limit at this order. More importantly, such expanded equations can also be obtained following Van Hoolst (1994), which takes a general set of linear eigenmodes as input to compute their nonlinear couplings and applies to generic compressible fluids (realistic stars and planets). This is advantageous over the affine model, which only self-consistently describes idealized homogeneous ellipsoids. Therefore, our study bridges the two different formulations of nonlinear hydrodynamics. The exact agreement between the two formulations at the three-wave order for homogeneous ellipsoids is presented in Appendix B.
@@
with $v_{m}=(q_{m}e^{im\phi})_{\phi=\phi_{q}}$. Treating $(-v_{0})\sim v_{2}\sim\xi$, we see that Eq. (10) takes the same form as the toy model of Eq. (1), and the nonlinear terms in $K_{m}$ tend to destabilize the system when $\xi>0$. Thus, instability is most likely to occur when $\xi$ reaches its maximum along the major axis. In the equilibrium limit (without mode resonances; following the convention of Yu et al. 2024), where the tidal bulge follows the companion, we have $\phi_{q}=\Phi$, matching the intuition that mass transfer first occurs at the L1 Lagrangian point. At the quadrupolar order, the L2 point ($\phi=\Phi+\pi$) destabilizes at the same time.
@@
The second correction comes from the centrifugal force whose detailed incorporation is presented in Appendix A. Since the force can be derived from a specific centrifugal potential $d\mathcal{R}/dM=-\omega^{2}(r^{2}-z^{2})/2$, its impact on the equations of motion can be obtained through a variational principle by considering the energy $\mathcal{R}$. When doing so, it is important to introduce an $l=0$ mode whose amplitude we denote as $q_{\rm tr}$ (a perturbation of the trace of $q_{ij}$ in Diener et al. 1995) and allow it to be coupled to the $q_{m}$ mode at the three-wave level to properly capture the correction at $\mathcal{O}(\epsilon^{2})$. The same result can also be obtained from the formulation of Van Hoolst (1994) by splitting $d\mathcal{R}/dM=(d\mathcal{R}/dM)_{l=0}+(d\mathcal{R}/dM)_{l=2}=-\omega^{2}r^{2}/3-\omega^{2}(x^{2}+y^{2}-2z^{2})/6$, and adding $(d\mathcal{R}/dM)_{l=0}$ to the background (monopolar) gravitational potential and $(d\mathcal{R}/dM)_{l=2}$ to the first-order, quadrupolar Eulerian potential perturbation (same as the tidal interaction potential). Note that our approach treats the centrifugal potential as a (nonlinear) drive of the modes defined by a non-rotating background, instead of redefining a new set of eigenmodes of a non-spherical background where the centrifugal perturbation is absorbed. In other words, the centrifugal potential is treated on an equal footing as the tidal potential.
@@
This splitting is based on the phase-space expansion of Schenk et al. (2002), and the Coriolis force is absorbed as a frequency shift (Lai, 2021; Yu et al., 2024; Yu and Lau, 2025),
@@
Higher-order corrections from the Coriolis force are ignored (but see, e.g., Dewberry and Lai 2022). The corresponding second-order equation can be obtained by differentiating $\dot{c}_{m,\pm}$ one more time and eliminating the first-order derivatives with the equations of motion, leading to
@@
Summarizing, the equations of motion we will solve in §II.1 and §II.2 are
@@
where $\bar{K}_{m}=K_{m}e^{im\phi_{q}}$ and $\bar{C}_{m}^{\rm(st)}=C_{m}^{\rm(st)}e^{im\phi_{q}}$. In the equilibrium limit (as considered in §II.1 and II.2) without free oscillations of the mode due to resonances, $\phi_{a}=\Phi_{C}$, and one just needs to replace $q_{m}$ with $v_{m}$ and drop the phases following $\epsilon$ in $K_{m}$ to obtain $\bar{K}_{m}$. Eq. (23) is real valued.
II.1 Consistency with previous analyses
Figure 1: Equilibrium sequences under the static limit. Curves in gray correspond to synchronized donors (with $p=1$), while those in yellow are for the non-rotating case ($p=0$). The solid and dashed lines are evaluated at the three-wave and four-wave orders, respectively. Vertical lines mark the locations where $\epsilon$ reaches the maxima. The classic results of Chandrasekhar (1963) are shown in the red horizontal lines.
First, we consider the case where $M$ moves in a circular orbit with synchronous rotation, so the left-hand sides of Eq. (20) or (23) vanish. Our goal is to establish the consistency between our formulation based on an order-by-order expansion with the classic analyses of Chandrasekhar (1963) under this static limit. We write the spin frequency $\omega$ as $\omega^{2}=p\epsilon$, with $p=1+M/M^{\prime}$ for synchronized rotation in a circular orbit. Note that our $p$ differs from the one in Chandrasekhar (1963) by 1, as $p=p^{\rm(Ch)}+1$. The superscript “(Ch)” will be used when we quote notations or results from Chandrasekhar (1963). In our analysis, we will focus on $M^{\prime}\gg M$ (as for planet migration or TDEs), so a synchronous donor has $p=1$.
Inspired by Chandrasekhar (1963), we note that the perturbed fluid can be described by an ellipsoid with $\bar{x}^{2}/a_{1}^{2}+\bar{y}^{2}/a_{2}^{2}+\bar{z}^{2}/a_{3}^{2}=1$ at the surface. The barred coordinates are aligned with the axes of the ellipsoid. In the limit with circular orbit and synchronous rotation, the barred frame coincides with the corotating frame. The axes are given by (Eqs. A5 and A31)
@@ II.1 Consistency with previous analyses
If there exists a $v_{2}$ that simultaneously solves the above two polynomials, then the resultant of the two polynomials must vanish (Weisstein, 2025). That is,
@@
In our calculation, we find that it is sufficient to capture the classic results of Chandrasekhar (1963) with percent-level accuracy using only three-wave interactions (solid lines; Eqs. 21 and 22). For the synchronous ($p=1$) case, we find $\epsilon_{\rm max}\simeq 0.0747$ at $\cos^{-1}a_{31}\simeq 64.8^{\circ}$, meaning the threshold separation is at $D^{(\rm th)}=(1/\epsilon_{\rm max})^{1/3}R_{t}\simeq 2.38R_{t}$, matching the classic Roche limit (indicated by the horizontal dotted lines) of $2.44R_{t}$ to $2.5\%$. A similar level of accuracy is also obtained for the non-rotating $p=0$ case, with $\epsilon_{\rm max}\simeq 0.110$, $(1/\epsilon_{\rm max})^{1/3}\simeq 2.09$, achieved at $\cos^{-1}a_{31}\simeq 65.4^{\circ}$.
The accuracy can be improved further by including the next order, four-wave interactions. The equations of motion at this order are derived in Appendix A, Eqs. (A33a) and (A33b). Still using Eq. (25) to eliminate $v_{0}$ in terms of $v_{2}$ and $a_{31}$, [^2] Strictly, at the four-wave order one should account for corrections to $a_{1,2,3}$ from the monopolar mode, $q_{\rm tr}$; Eqs. (A5) and (A31). This introduces nonlinear corrections to Eq. (25), which we ignore here. we again obtain two polynomial equations of $v_{2}$,
@@
by setting $\ddot{q}_{m}=0$, with $m=2$ and 0. The resultant (now functions of $W_{mn}$) needs to vanish as before for the two equations to hold simultaneously. We can thus find the equilibrium sequences at the four-wave order, shown in the dashed lines in Fig. 1. The result now matches those from Chandrasekhar (1963) almost exactly. For the $p=1$ case shown in gray, we have $\epsilon_{\rm max}\simeq 0.0687$ at $\cos^{-1}a_{31}=61.1^{\circ}$, corresponding to a Roche limit of $D^{(\rm th)}=(1/\epsilon_{\rm max})^{1/3}R_{t}=2.44R_{t}$.
@@
While the main result of this Subsection is to reproduce the classic ones by Chandrasekhar (1963) on homogeneous ellipsoids, it is significant for two reasons. First, it validates our formulation based on which $\ddot{q}_{m}\neq 0$ corrections will be introduced. Second, it bridges the analyses of Chandrasekhar (1963) on idealized ellipsoids and Van Hoolst (1994) on general stars and planets (see also Weinberg et al. 2012; Weinberg 2016; Yu et al. 2023). In general, exact nonlinear expressions are challenging to find, but an order-by-order expansion as in Van Hoolst (1994) is straightforward to compute (though lengthy). Our result provides a framework for how one can start from Van Hoolst (1994) to determine the Roche limit of a realistic star/planet theoretically.
II.2 Finite-frequency corrections in asynchronous, low eccentricity systems
@@
In this Subsection, we will extend the stability analysis to analytically consider finite-frequency corrections due to the equilibrium tide (following the convention of Yu et al. 2024), where the perturbation varies with the instantaneous tidal forcing. In other words, the tidal bulge still points towards the companion. The component that varies at a mode’s natural frequency and depends on the excitation history of the mode (what we call the dynamical tide) will be deferred to the numerical investigation in §II.3 and ignored in the analysis here. This means that the prediction derived here will be accurate for binaries with low eccentricity $e\lesssim 0.3$. For more eccentric binaries, the dynamical tides can exhibit complicated, often diffusive, evolutions that may trigger instabilities at separations greater than the prediction of this Subsection by tens of percent.
@@
To proceed, we write $q_{m}=(1+z_{m})q_{m}^{(0)}$ where $q_{m}^{(0)}\simeq I_{m}\epsilon e^{-im\Phi_{C}}$ represents an approximate solution of the perturbation, with $I_{2}=3/4$ and $I_{0}=-\sqrt{6}/4$ (Eqs. 21 and 22). We treat $z_{m}$ as a low-frequency variable (whose time derivatives will be ignored) that we will use to determine the equilibrium structure of the donor. Note Eq. (23) still holds as long as we incorporate the time derivatives on the left-hand side due to $\ddot{q}_{m}\simeq\ddot{q}_{m}^{(0)}$. In essence, this model corresponds to the toy model described in Eq. (6).
@@
The procedure for finding the Roche limit follows identically those described in §II.1. Choosing again a linear relation between $v_{0}$ and $v_{2}$ as in Eq. (25), Eq. (23) now becomes two quadratic equations of $z_{2}$:
@@
For practical usefulness, we show in Fig. 2 the threshold pericenter separation, $D_{p}^{(\rm th)}/R_{t}=(1/\epsilon_{\rm max})^{1/3}$, below which hydrodynamical instability can occur. Different colors represent different spin rates relative to the pseudo-synchronization values $\omega_{\rm ps}$ in Hut (1981),
@@
We pause here to discuss further the meaning of Fig. 2. As the coefficients in Eq. (30) are evaluated at the pericenter, Fig. 2 should be interpreted as the instantaneous instability threshold at the pericenter. Therefore, a donor $M$ marginally unstable at the pericenter may become stable once it moves away from the pericenter and $\epsilon$ (defined by the instantaneous separation) drops below the instability threshold. As a result, the growth of hydrodynamical instability during a single pericenter passage may not disrupt $M$ sufficiently to cause mass loss. Nonetheless, as long as the fluid dissipation timescale is long compared to the orbital period (which is very well satisfied for the large-scale f-mode when 1e>0.011-e>0.01 under the linear theory assuming damping due to turbulent convection; Goldreich and Nicholson 1977; Wu 2005), the growth of instability at every pericenter passage will accumulate over multiple orbits, eventually disrupting $M$ and triggering mass loss (see also numerical examples to be presented in the next Subsection). For systems with low eccentricities $e\lesssim 0.3$, Fig. 2 provides a decent approximation of the onset of mass loss as a runaway in mode amplitude occurs within a few (<10<10) orbits. For systems with higher eccentricities, the amount of growth per passage is smaller, given the smaller fraction of time $M$ spends near the pericenter ($2\pi/(\dot{\Phi}_{p}P_{\rm orb})\propto\sqrt{(1-e)^{3}/(1+e)}$). This requires more orbits for the perturbation to build up until a runaway eventually occurs. However, the dynamical tide component can also be excited in those systems, significantly lowering the threshold $\epsilon$ for a stable configuration (see Eq. 6 and note the dynamical tide plays the role of $\xi_{\rm dyn}$ in the toy model). In the weak damping case, mass loss from $M$ can occur at $D_{p}$ significantly greater than the prediction of Fig. 2 for highly eccentric systems.
II.3 Dynamical tide induced instability and diffusive tide